A ball is thrown horizontality from a height of 20m and hits the ground with a speed that is three times it’s initial speed. What is the initial speed?


1
Expert's answer
2019-12-16T11:11:07-0500

We have:


"v_f^2=v^2+2gh"

"(3v)^2=v^2+2gh\to 8v^2=2gh"

"8v^2=2(9.8)(20)"

"v=7\frac{m}{s}"